$SPAD/src/input stbl.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool stbl.output
)set message test on
)set message auto off
)clear all
--S 1 of 7
t: SparseTable(Integer, String, "Try again!") := table()
--R
--R
--R   (1)  table()
--R                                 Type: SparseTable(Integer,String,Try again!)
--E 1
--S 2 of 7
t.3 := "Number three"
--R
--R
--R   (2)  "Number three"
--R                                                                 Type: String
--E 2
--S 3 of 7
t.4 := "Number four"
--R
--R
--R   (3)  "Number four"
--R                                                                 Type: String
--E 3
--S 4 of 7
t.3
--R
--R
--R   (4)  "Number three"
--R                                                                 Type: String
--E 4
--S 5 of 7
t.2
--R

--R
--R   (5)  "Try again!"
--R                                                                 Type: String
--E 5
--S 6 of 7
keys t
--R
--R
--R   (6)  [4,3]
--R                                                          Type: List(Integer)
--E 6
--S 7 of 7
entries t
--R
--R
--R   (7)  ["Number four","Number three"]
--R                                                           Type: List(String)
--E 7
)spool
)lisp (bye)

———-

References^{[1] nothing}

5